| Operation Summary for April 23, 2012 - April 25, 2012 |
|
Monday April 23 The day shift began with Operations monitoring the Linac, Booster, and Main Injector (MI), and with the NuMI, SeaQuest, and the Meson FTBFT-1015 experiments taking beam. MiniBooNE began its shutdown on the midnight shift. During the day shift, a Muon Ring person began about a one-hour study at 10 AM. At 10:26 AM, the NuMI beam permit tripped off due to a power supply (E:V118) drifting low; everything reset. At 11:56 AM, the NuMI beam permit tripped off due to power supply (E:HV101) tripping off: everything reset. Operations established beam to MTA at 12:31 PM. Operations prepared the Neutrino Area and SeaQuest for access at 1:02 PM. A MI person began a MI-52 Septa beam scan at 1:06 PM. At 3:13 PM, SeaQuest resumed taking beam. MI-52 Septa studies finished at 3:28 PM. Muon Ring studies finished at 3:47 PM. During the evening shift, MTA turned off at 10:51 PM. Tuesday April 24 The midnight shift began with Operations sending beam to NuMI, SeaQuest, and T-1015. During the midnight shift, at 1:08 AM, the NuMI beam permit tripped off due to power supply (E:HV101) tripping off: everything reset. During the day shift, at 11:02 AM, the NuMI beam permit tripped off due to power supply (E:V108) tripping off on “LCW temp hot”: everything reset. MTA turned on at 11:26 AM. Muon Ring personnel began studies at 11:48 AM. Muon Ring studies finished at 3:45 PM. During the evening shift, Operations delivered beam to users. Wednesday April 25 The midnight shift began with Operations sending beam to NuMI, SeaQuest, and T-1015. During the midnight shift, MTA turned off at 1:42 AM. MTA resumed taking beam at 7:00 AM. Operations Coordinator (OpsCo) Plans for Today Access the Muon transport enclosure at 10:00 AM. After the access deliver beam to users. |
|
More Information Comments and Suggestions |
| last modified 04/25/2012 email Fermilab |
| Security, Privacy, Legal | Use of Cookies | |